The following is the first quote:

>>  TEX is an accessible and effective tool for typesetting
>> written documents. Users at all levels of experience, however,
>> sometimes require help not readily available.
>> 1. This is the first item. This is the first item. This is the
>>  first item. This is the first item.
>> 2. This is the second item. This is the second item. This
>>  is the second item. This is the second item.

The following is the second quote:

||  1. Calculate 2^{3^{4^{5}}} .
||  2. Calculate ((2^{3})^{4})^{5}.